OpenStudy (anonymous):

Solve. 2x2 + 16x + 50 = 0 Answer

OpenStudy (anonymous):

This one use quadratic formula. a=2, b=16, c=50 (-16 + or - sqrt(16^2 - 4(2)(50)))/(2*2)

OpenStudy (anonymous):

(-16 + or - sqrt(-656))/(4)

OpenStudy (anonymous):

(-16 + or - 4sqrt(-41))/(4) then factor out 4 4(-4 + or - sqrt(-41))/(4) and then x = -4 + sqrt(-41) x = -4 - sqrt(-41)

jimthompson5910 (jim_thompson5910):

16^2 - 4(2)(50) = 256 - 400 = -144 So you should have (-16 + or - sqrt(-144))/(4)

OpenStudy (anonymous):

damn I did -(16^2) for some reason Jim is right^

jimthompson5910 (jim_thompson5910):

everything else is right though, so good job

jimthompson5910 (jim_thompson5910):

well I mean everything before it of course

OpenStudy (anonymous):

Thanks =) I need to pay closer attention though. haha yes.

jimthompson5910 (jim_thompson5910):

and the ideas are still valid, so the OP still gets a good idea what to do

jimthompson5910 (jim_thompson5910):

yw